Expand and simplify.
3
E(1/(n^2))
n=1

Answer by oscargut(2103)

= 1/1^2 + 1/2^2 +`1/3^2
= 1+1/4+1/9
= 49/36
If you have more questions i can see them if you contact me at: mthman@gmail.com
Thanks